OpenStudy (ayeeeitskateee):

How are the functions y=x and y=x - 3 related? How are their graphs related? 1.) Each output for y=x-3 is 3 more than the corresponding output for y=x. the graph for y=x-3 is translated down 3 units. 2.) Each output for y=x-3 is 3 more than the corresponding output for y=x. The graph of y=x-3 is the graph of y=x translated up 3 units. 3.) Each output for y=x-3 is 3 less than the corresponding output for y=x. The graph for y=x-3 is 3 less then the corresponding output for y=x. will medal/fan.

OpenStudy (ayeeeitskateee):

4.) Each output for y=x-3 is 3 less than the corresponding output for y=x. The graph for y=x-3 is the graph of y=x translated up 3 units.